# AuscultaBase: A Foundational Step Towards AI-Powered Body Sound Diagnostics

### v1

@@ Abstract
Auscultation of internal body sounds is essential for diagnosing a range of health conditions, yet its effectiveness is often limited by clinicians’ expertise and the acoustic constraints of human hearing, restricting its use across various clinical scenarios. To address these challenges, we introduce AuscultaBase, a foundational framework aimed at advancing body sound diagnostics through innovative data integration and contrastive learning techniques. Our contributions include the following: First, we compile AuscultaBase-Corpus, a large-scale, multi-source body sound database encompassing 11 datasets with 40,317 audio recordings and totaling 322.4 hours of heart, lung, and bowel sounds. Second, we develop AuscultaBase-Model, a foundational diagnostic model for body sounds, utilizing contrastive learning on the compiled corpus. Third, we establish AuscultaBase-Bench, a comprehensive benchmark containing 16 sub-tasks, assessing the performance of various open-source acoustic pre-trained models. Evaluation results indicate that our model outperforms all other open-source models in 12 out of 16 tasks, demonstrating the efficacy of our approach in advancing diagnostic capabilities for body sound analysis.
@@ 1 Introduction
Auscultation, the clinical skill of listening to internal body sounds, is essential for diagnosing a variety of health conditions. By analyzing heart sounds, respiratory sounds, and bowel sounds, clinicians can derive critical insights into a patient’s physiological state, enabling the monitoring of conditions such as valvular heart disease, congenital heart defects, pneumonia, asthma, gastroenteritis, and intestinal obstruction. However, the efficacy of auscultation is largely contingent upon the clinician’s expertise and the acoustic range of the human ear, which limits its application in various clinical scenarios (Mangione & Nieman, 1997). Recently, the integration of artificial intelligence algorithms for sound analysis has emerged as a promising approach to enhance diagnostic accuracy (Chen et al., 2023; Kong et al., 2020; Oletic & Bilas, 2017; Zhang et al., 2024; Sitaula et al., 2022).
Despite the progress made with existing methodologies, several challenges persist. First, while some approaches utilize general acoustic techniques to extract audio features from specialized stethoscope recordings, they may not adequately capture the pathological variations in body sounds (Chen et al., 2023; Guo et al., 2023; Oletic & Bilas, 2017; Jung et al., 2021). With advancements in unsupervised training, neural networks have been shown to effectively learn high-quality representations directly from audio data (Baevski et al., 2020; Huang et al., 2022; Hsu et al., 2021b; Wu et al., 2023). Some methods have proposed the use of these pretrained audio models for feature extraction for body sound (Koike et al., 2020; Panah et al., 2023). However, existing pretrained models are typically trained on general audio datasets like AudioSet (Gemmeke et al., 2017) and Librispeech (Panayotov et al., 2015) which contain sounds—such as music and speech—that significantly differ from those collected by stethoscopes, leading to substantial domain gaps that render them unsuitable for auscultation feature extraction. Currently, there are publicly available datasets for the three types of body sounds and the details are shown in Tabel 1. Zhang et al. (2024) utilized publicly available respiratory sounds captured by stethoscopes alongside cough sounds recorded by microphones to develop respiratory sound representations. Baur et al. (2024) identified and utilized cough sounds, respiratory sounds, and laughter collected from YouTube to construct health-related acoustic representations. However, these methodologies are not specifically designed for auscultation and may not effectively adapt to the diverse range of body sounds captured by stethoscopes.
To address these critical gaps, we introduce AuscultaBase, a foundational framework designed to transform body sound diagnostics through advanced data integration and contrastive learning techniques. AuscultaBase offers a comprehensive solution that not only supports the diagnostic capabilities of primary care providers but also enables scalable, high-quality auscultation analysis, particularly for clinicians operating in high-volume, low-resource environments. Our framework is built upon three core pillars:
@@
  • AuscultaBase-Corpus: A Large-Scale, Multi-Source Body Sound Database. This extensive corpus combines 10 publicly accessible datasets and one private dataset, covering a diverse array of body sounds—heart, respiratory, and gastrointestinal—for a total of 322.4 hours of training recordings. The AuscultaBase-Corpus provides a rich, representative data foundation to train and evaluate diagnostic models across multiple auscultation systems.
  • AuscultaBase-Model: A Foundational Diagnostic Model for Body Sounds. We developed a robust foundational model tailored for body sound diagnostics using contrastive learning techniques on the AuscultaBase-Corpus. This model captures nuanced acoustic patterns, showing significant improvements in diagnostic accuracy over conventional acoustic models. It offers a unified approach to analyzing different body sounds, setting a new standard for accuracy and reliability in auscultation.

### v2

# A Generalist Audio Foundation Model for Comprehensive Body Sound Auscultation

@@ Abstract
Accurate and efficient auscultation-based diagnostics are vital for early disease detection, especially in resource-limited settings where specialized clinical expertise is scarce. Traditional auscultation, which heavily depends on clinician experience, suffers from significant inter-observer variability, while existing AI models often falter due to the limitations of non-representative training data. In this study, we introduce AuscultaBase, a novel AI-driven diagnostic framework that harnesses self-supervised and contrastive learning techniques alongside large-scale, multi-source data integration to advance body sound analysis. By generating robust feature representations, AuscultaBase markedly enhances performance in abnormality detection, disease classification, and activity recognition tasks. Comprehensive evaluations on our newly established benchmark, AuscultaBench, demonstrate that AuscultaBase consistently outperforms state-of-the-art methods across key performance metrics, underscoring its potential as a scalable and cost-effective tool for clinical screening and early disease intervention. The code and model checkpoint has been released in https://github.com/applewpj/AuscultaBase.
@@ 1 Introduction
Auscultation, the clinical practice of listening to internal body sounds, is a cornerstone of non-invasive diagnosis for a range of health conditions. By analyzing heart, respiratory, and bowel sounds, clinicians gain critical insights into a patient’s physiological state, enabling the detection and monitoring of disorders such as valvular heart disease, congenital heart defects, pneumonia, asthma, gastroenteritis, and intestinal obstruction. However, the accuracy and consistency of auscultation remain heavily reliant on clinician expertise and the limited acoustic sensitivity of the human ear, which can result in variability and missed diagnoses (Mangione & Nieman, 1997). The growing need for standardized, efficient, and widely accessible diagnostic tools has driven research into artificial intelligence (AI)-based auscultation analysis, which offers the potential to enhance diagnostic accuracy and scalability (Chen et al., 2023; Kong et al., 2020; Oletic & Bilas, 2017; Zhang et al., 2024; Sitaula et al., 2022).
Despite recent progress, several fundamental challenges remain unaddressed. Existing methods predominantly rely on general acoustic techniques that are not optimized for auscultatory signals, limiting their ability to capture the pathological variations in body sounds (Chen et al., 2023; Guo et al., 2023; Oletic & Bilas, 2017; Jung et al., 2021). While unsupervised neural network training has demonstrated effectiveness in learning high-quality audio representations (Baevski et al., 2020; Huang et al., 2022; Hsu et al., 2021b; Wu et al., 2023), pre-trained audio models are often developed using datasets such as AudioSet (Gemmeke et al., 2017) and LibriSpeech (Panayotov et al., 2015), which primarily contain music and speech rather than medically relevant auscultatory sounds. This domain gap diminishes their applicability for body sound analysis. Furthermore, publicly available datasets currently span only a fraction of the diversity observed in clinical auscultation, further constraining the generalization ability of AI-driven models.
Given the critical need for rapid and cost-effective screening solutions, especially in resource-limited settings, we propose a comprehensive framework designed to advance AI-assisted auscultation. Our approach aims to enhance the diagnostic capabilities of primary care providers while enabling scalable, high-quality body sound analysis. The framework is structured around three core components:
@@
  • AuscultaCorpus: A large-scale, multi-source body sound database. This corpus constitutes an extensive repository of body sound recordings, integrating ten publicly available datasets with one proprietary dataset. This amalgamation encompasses a diverse spectrum of auscultatory sounds, including cardiac, respiratory, and gastrointestinal recordings, totaling 322.4 hours of training data. The corpus offers a comprehensive and representative foundation for the development and evaluation of diagnostic models across various auscultation systems, thereby facilitating advancements in automated diagnostic methodologies.
  • AuscultaBase: A foundational diagnostic model for body sounds. AuscultaBase is developed as a robust diagnostic framework specifically designed for body sound analysis. Built on contrastive learning techniques applied to the AuscultaCorpus, the model effectively captures intricate acoustic patterns inherent in physiological sounds. Comparative evaluations demonstrate significant enhancements in diagnostic accuracy relative to traditional acoustic models. This unified approach enables the simultaneous analysis of diverse body sounds, thereby establishing a new benchmark for precision and reliability in auscultation-based diagnostics.